# Aliasing from Galactic Plane Setting in Widefield Radio Interferometry

### v1

@@ 5.1 Extragalactic Catalogs
Figure 3: The reconstructed 2D (left) and 1D (right) power spectra of two hours of simulated Galactic plane data with the MWA. There are no other sources of power or noise. Greyed regions in the 2D power spectra are excluded from the cylindrical averaging which creates the 1D power spectra. Depending on window leakage and flagged channels, the Galactic plane power (purple) can be greater than the EoR signal itself (orange). Not accounting for missing/flagged data is particularly harmful (solid purple), given that the Galactic plane along the horizon line is coupled to the missing/flagged modes. This results in the characteristic double peaks seen in the 1D power spectra.
@@
The EoR power spectrum (PS) measurement relies on the natural separation of foregrounds from the 21-cm signal in Fourier space. Noise-like signals which do not decorrelate with time could potentially preclude the measurement.
@@
PS analysis typically generates either a measured PS (i.e. delay or gridded delay, Parsons et al. 2012) or a reconstructed PS (i.e. {$k_{\perp}$, $k_{||}$}, Morales & Hewitt 2004). Thyagarajan et al. (2015) first measured the implications of widefield foregrounds, including the Galactic plane, on measured PS analyses. In this work, we demonstrate the effects of the Galactic plane setting with simulations in a reconstructed PS analysis.
@@
Figure 3 shows the 2D PS of simulations of the Galactic plane through the MWA for a 2-hr data set centered on field EoR0 (left). Foregrounds are generally confined to the region below the solid black line, as is characteristic with the expected chromaticity of the instrument (Datta et al., 2010; Morales et al., 2012; Parsons et al., 2012; Trott et al., 2012; Vedantham et al., 2012; Hazelton et al., 2013; Pober et al., 2013; Thyagarajan et al., 2013; Liu et al., 2014). The location of the foregrounds on the sky will affect the associated location of power in the PS space. As the foregrounds move away from the main Field-of-View towards the horizon, they will move as a $k_{\perp}\propto k_{||}$ line, where the intercept increases with distance from phase center (i.e. Morales et al. 2012; Pober et al. 2016). This line is shown in Figure 3 (left) as a solid black line for foregrounds at the horizon.
@@
Our simulations model the data from the MWA. As such, we include regular flagging in frequency due to aliased channels from the polyphase filter bank, which are then averaged in frequency. The foregrounds are coupled to the channels with incomplete, missing data. This manifests as contaminated lines in $k_{||}$, three of which can be seen in Figure 3. The original locations of foregrounds are reflected in the contaminated $k_{||}$ modes, creating a double-pronged feature with the Galactic plane which is reminiscent of the ”pitchfork” effect (see Thyagarajan et al. 2015).
When cylindrically averaged to a 1D PS (right, solid purple line), these double-pronged features contaminate enough modes to preclude the EoR measurement (right, orange line). However, the effects of missing or incomplete channels can be mitigated by employing extensive removal techniques or full covariance weighting (i.e. Offringa et al. 2019). We also show a 1D PS of unflagged Galactic plane simulations (right, dashed purple line), which show that measurements of the 21.
[…]
@@ 6 Conclusion
Widefield radio interferometry can be affected by aliasing inherent to the measurement itself from the bright, Galactic plane setting over the opaque horizon. We show simulations of this effect in great detail for the first time using the diffuse catalog from Kriele et al. (2022) as input into WODEN, a GPU-enabled analytic simulator (Line, 2022), to create visibilities that are imaged with FHD (Sullivan et al., 2012; Barry et al., 2019a). These simulations, when subtracted from data, show improved image quality in just 2 min of observation.
@@
The noise statistics of a real data image are improved with the removal of the Galactic plane and its aliasing. The resulting MWA image has less variance, skewness, and kurtosis. For future measurements of higher-order moments of the EoR (Watkinson & Pritchard, 2015; Kittiwisit et al., 2022), like the skew spectrum (Ma & Peng, 2023; Cook et al., in review), this could have a huge potential impact if done in image space.
@@
    For extragalactic cataloging, a small bias is introduced in the RMS maps of the mosaic image, which in turn propagates to the number of sources recovered and their integrated flux. This is, in general, small enough to not be of concern when a range of LST are used to create the mosaic image.
@@
    For 21 cm EoR power spectra detection, there is quite a substantial effect near the horizon line in $\{k_{\perp},k_{||}\}$ as expected. If there is regular flagging in frequency space, as is the case for the MWA, this shape will be reflected, creating mirrors of the horizon line that occupy a large number of $k$-modes. Reconstructed PS analyses which take an avoidance strategy will not be able to measure the EoR if this effect is not mitigated.
Our simulations and real data analysis show that the setting of the Galactic plane can introduce aliasing into both Fourier-space and image-space statistical metrics. Given the brightness of the Galactic plane, beam attenuation of 0.1% is not enough to remove this effect from precision science goals.

### v2

@@ 5.2 The 21 cm EoR Power Spectrum
Figure 3: The reconstructed 2D (left) and 1D (right) power spectra (PS) of two hours of simulated Galactic plane data with the MWA. There are no other sources of power or noise. The contour on the 2D PS highlights bins used to generate 1D PS. The simulated Galactic plane power (solid purple) is just below the theoretical EoR signal (orange line with 95% confidence region). Basic frequency-dependent calibration errors from an extragalactic calibration cause excess contamination (dashed purple). Not accounting for missing/flagged data is particularly harmful (solid purple), given that the Galactic plane along the horizon line is coupled to the missing/flagged modes.
@@
The EoR power spectrum (PS) measurement relies on the natural separation of foregrounds from the 21 cm signal in Fourier space. Noise-like signals which do not decorrelate with time could potentially preclude the measurement.
@@
PS analysis typically generates either a measured PS (e.g., delay or gridded delay, Parsons et al. 2012) or a reconstructed PS (e.g., {$k_{\perp}$, $k_{||}$}, Morales & Hewitt 2004). Thyagarajan et al. (2015) first measured the implications of widefield foregrounds, including the Galactic plane, on measured PS analyses. In this work, we demonstrate the effects of the Galactic plane setting with simulations in a reconstructed PS analysis.
@@
The left panel of Figure 3 shows the 2D PS of simulations of the Galactic plane through the MWA for a total of 2 hr centered on field EoR0. The location of the foregrounds on the sky will affect the associated location of power in the PS space. As the foregrounds move away from the main FoV towards the horizon, they will move as a $k_{\perp}\propto k_{||}$ line, where the intercept increases with distance from phase center (e.g., Morales et al. 2012; Pober et al. 2016). This line is shown in Figure 3 (left) as a solid black line for foregrounds at the horizon. Foregrounds are generally confined to the region below the solid black line, as is characteristic with the expected chromaticity of the instrument (Datta et al., 2010; Morales et al., 2012; Parsons et al., 2012; Trott et al., 2012; Vedantham et al., 2012; Hazelton et al., 2013; Pober et al., 2013; Thyagarajan et al., 2013; Liu et al., 2014).
@@
Our simulations model the data from the MWA. As such, we include regular flagging in frequency due to aliased channels from the polyphase filter bank, which are then averaged in frequency. The foregrounds are coupled to the channels with incomplete, missing data. This manifests as contaminated lines in $k_{||}$. The original locations of foregrounds are reflected in the contaminated $k_{||}$ modes, creating double-pronged features with the Galactic plane which can be seen in the left panel of Figure 3. This is reminiscent of the “pitchfork” effect (see Thyagarajan et al. 2015).
We cylindrically average the 2D PS within the contoured region of the left panel of Figure 3 to investigate how the contamination propagates to the 1D PS. The orange line and shaded area in the right panel shows the expected EoR measurement with 95% confidence (Barry et al., 2019b). The double-pronged features of the missing channels in the simulated Galactic plane (dotted purple) preclude the EoR signal in all modes in 1D space.
If there were no missing or incomplete channels, then the Galactic plane simulation is below the EoR signal (solid purple) [^9] Remaining power above $k\approx 0.4\,\textit{h}$ Mpc-1 is most likely due to unrelated analysis systematics in the beam model during the generation of the model visibilities.. The effects of missing or incomplete channels can be mitigated by employing extensive removal techniques or full covariance weighting (e.g., Offringa et al. 2019).
We also show the consequences of excluding the Galactic plane in basic calibration in the 1D PS (dashed purple). We use the per-frequency gain solutions from a linear least-squares solver between (1) visibilities including extragalactic sources and the Galactic plane, and (2) visibilities from only extragalactic sources. This type of calibration is generally the starting point for most modern calibration techniques.
We exclude baselines less than 50 $\lambda$ from the calibration (Patil et al., 2016), a common technique to avoid model dependence on diffuse emission. Nevertheless, the Galactic plane aliasing still affects the calibration such that there is at least an order of magnitude contamination in modes greater than 0.4 h Mpc-1. This is due to a significant amount of power from this diffuse structure being greater than 50 $\lambda$ (Byrne et al., 2022), and the extreme amount of precision required in this basic, per-frequency calibration formalism (Barry et al., 2016). Most calibration procedures are now much more advanced – for example, we use an auto-correlation formalism in calibrating our data in Section 4.
Figure 3 focuses solely on simulations. However, we can also briefly investigate the effects on the EoR PS using real data. Using 2 hr of data centered on EoR0 as an example, we generate limits on the 21 cm signal. We use the same data selection and binning techniques as Barry et al. (2019b) for simplicity. The data was processed in two ways—with and without subtracting the Galactic plane. The calibration and baseline selection was kept the same to create a fair comparison. When the Galactic plane is removed in addition, there is a reduction in the upper limit by 0.4% in the N–S polarization at 0.2 h Mpc-1, which was the lowest $k$-mode in Barry et al. (2019b). This reduction, whilst seemingly small, is still over 20x larger than the 21 cm signal itself. Future work with real MWA data centered on EoR0 will investigate differences in data on a larger scale.
@@ 6 Conclusion
Widefield radio interferometry can be affected by aliasing inherent to the measurement itself from the bright, Galactic plane setting over the opaque horizon. We show simulations of this effect in great detail for the first time using the diffuse catalog from Kriele et al. (2022) as input into WODEN, a GPU-enabled analytic simulator (Line, 2022), to create visibilities that are imaged with FHD (Sullivan et al., 2012; Barry et al., 2019a). These simulations, when subtracted from data, show improved image quality in just two minutes of observation.
@@
The noise statistics of a real data image are improved with the removal of the Galactic plane and its aliasing. The resulting MWA image has less variance, skewness, and kurtosis. For future measurements of higher-order moments of the EoR (Watkinson & Pritchard, 2015; Kittiwisit et al., 2022), like the skew spectrum (Ma & Peng, 2023; Cook et al., 2024), this could have a huge potential impact if done in image space.
@@
    For extragalactic cataloging, the Galactic plane aliasing can be subsumed into the calculated rms maps of the mosaic image. This artificially boosts the subtracted rms, changing the amount of detected faint sources and their flux density. This is, in general, small enough to not be of concern when a range of LST are used to create the mosaic image. For example, with 2 hr of MWA simulation, we see a difference of less than 0.5% recovered sources, and changes in flux densities of recovered sources of less than 200 mJy.
@@
    For 21 cm EoR power spectra detection, there is quite a substantial effect near the horizon line in $\{k_{\perp},k_{||}\}$ as expected. If there is regular flagging in frequency space, as is the case for the MWA, this shape will be reflected, creating mirrors of the horizon line that occupy a large number of $k$ modes. Basic per-frequency calibration formalisms will be affected by the Galactic plane aliasing, even if short baselines are excluded. Reconstructed PS analyses must consider the Galactic plane if it is setting, or at least employ a variety of techniques to remove its effects in flagging, calibration, and foreground power.
Our simulations and real data analysis show that the setting of the Galactic plane can introduce aliasing into both Fourier-space and image-space statistical metrics. Given the brightness of the Galactic plane, beam attenuation of 0.1% is not enough to remove this effect from some precision science goals.